# Structure and Dynamics of Magneto-Inertial, Differentially Rotating Laboratory Plasmas

### v1

@@ 5 Conclusions
We have presented a detailed characterization of the structure and dynamics of rotating plasmas driven on the RPX pulsed-power platform. This report complements a previous publication by Valenzuela-Villaseca et al. (2023a), as follows:
@@
    We have derived a self-similar solution describing the radial growth and velocity evolution of the rotating plasma as it accretes mass and angular momentum from the ablation flows, which also provide radial confinement through ram pressure. The model shows that the rotation velocity gradually decreases as the plasma moment of inertia increases. However, this decay saturates at an estimated the velocity of $u_{\theta}\sim 20$ km/s, which is in agreement with direct TS measurements. A straight forward application of this result is to couple this self-similar solution with the calculated magnetic field components presented earlier in the paper, to easily design alternative RPX configurations and generate plasmas with a given radius and rotation velocity, without the need for computationally expensive simulations.
@@ Acknowledgements
This work was supported in part by NNSA under DOE Cooperative Agreement No DE-SC0020434 and DE-NA0003764. Vicente Valenzuela-Villaseca was funded by the Imperial College President’s PhD Scholarships and the Royal Astronomical Society. We are grateful to the Mechanical Workshop at the Imperial College Department of Physics, and in particular to David Williams, for hardware fabrication. We thank Philipp Kempski for helpful conversations about the equations presented in the discussion. We thank Thomas Varnish for writing some of the codes used in the analysis.

### v2

@@ 5 Conclusions
We have presented a detailed characterization of the structure and dynamics of rotating plasmas driven on the RPX pulsed-power platform. This report complements a previous publication by Valenzuela-Villaseca et al. (2023), as follows:
@@
    We have derived a self-similar solution describing the radial growth and velocity evolution of a thin, rotating plasma shell. The rotating plasma grows and slows down as it accretes mass and angular momentum from the ablation flows, which also provide radial confinement through ram pressure. The model shows that the rotation velocity gradually decreases as the plasma moment of inertia increases. However, this decay saturates at an estimated the velocity of $u_{\theta}\sim 20$ km/s, which is in agreement with direct TS measurements. We find that both the outer boundary of the plasma in the experiment and measured rotation velocity are consistent the self-similar trajectory. A straight forward application of this result is to couple this self-similar solution with the calculated magnetic field components presented earlier in the paper, to easily design alternative RPX configurations and generate plasmas with a given radius and rotation velocity, without the need for computationally expensive simulations.
@@ Acknowledgements
We are grateful to the Mechanical Workshop at the Imperial College Department of Physics, and in particular to David Williams, for hardware fabrication. We thank Philipp K.-S. Kempski for fruitful conversations about the equations presented in the discussion. We thank Thomas Varnish for writing some of the codes used in the analysis. Vicente Valenzuela-Villaseca is currently at the Department of Astrophysical Sciences, Princeton University, Princeton, New Jersey, USA. Francisco Suzuki-Vidal is currently at First Light Fusion Ltd., Yarnton, UK. Jack W. D. Halliday is currently with at the Department of Physics, Clarendon Laboratory, University of Oxford, Oxford, UK. Danny R. Russell is currently at the Technische Universitaet Muenchen, Forschungs Neutronenquelle Heinz Maier-Leibnitz, Munich, Germany.
Funding
This work was supported in part by NNSA under DOE Cooperative Agreement No DE-SC0020434 and DE-NA0003764. The work of Vicente Valenzuela-Villaseca was supported by the Imperial College President’s PhD Scholarships and the Royal Astronomical Society.
Declaration of Interests
The authors report no conflict of interest.